Team:Calgary/11 June 2010


Friday June 11, 2010

Alex, Patrick, and Raida's R0040-E0430 construct.

Dev, Jeremy, and Chris

Today, we spent most of the day preparing tubes of competent cells which could be used in future transformations from an original stock of Top10. As well, the overnight cultures of the plasmid transfers of E0032 and E0040 were Miniprepped using GenElute Plasmid Preparation procedure so the plasmids could be constructed on Monday. The construction was not grown over the weekend so as not to overgrow the plates. In the afternoon, we had a group meeting with Dr. Cairine Logan (Our teacher sponsor) and Deirdre Lobb (Our laboratory technician) where we discussed the various aspects of the project, gave reports on what we had accomplished so far, diagrammed the circuits so we knew exactly what was going on, and made plans for the following week. One idea that came up during the meeting was the use of a Transcriptional Terminator to be the negative control so the systems would not be expressed. Paul (One of our advisors) presented a possible modelling project that we could incorporate using a Neural Network system in MatLab. More information to follow on this later.

Alex, Patrick, and Raida

It was only this morning did we realize that it had probably not been a good idea to have both the vector and insert be ampicillin-resistant, as there is no selection pressure for the parts that did not work. We will have to do a plasmid switch for one of the parts (E0430) to pSB1K2, but we will leave that to Monday.

On another note, we selected four of what appeared to be the yellowest-looking colonies from the plate and restreaked them. Hopefully they will be yellower over the weekend.


Today, I ran another PCR of the construction of B0034-E1010 as no lines showed up on the agarose gel yesterday. No lines were present once again when gel electrophoresis was done. Yesterday's overnight cultures of E1010 and B0034-E1010 construct were miniprepped. I am unsure as to why the PCRs and gel electrophoresis are not working. Some possible reasons could be that the DNA is too short (expected range is approximately 700 bp), the gel is running for too long (it ran for 40-50 minmutes), or no positive control was present. On the next attempt, I will run a positive control for the parts of interest.